# Your Classifier Can Do More: Towards Bridging the Gaps in Classification, Robustness, and Generation

### v1

@@ Abstract
Joint Energy-based Models (JEMs), a class of hybrid generative-discriminative models, are well known for their ability to achieve both high classification accuracy and generative capability within a single model. However, their robustness still lags significantly behind the classifiers based adversarial training (AT). Conversely, while AT is currently the most effective approach to improving the classifier’s robustness, it typically sacrifices accuracy on clean data and lacks generative capability. The triple trade-off between classification accuracy, generative capability and robustness, raises a natural question: Can a single model simultaneously achieve high classification accuracy, adversarial robustness, and generative performance? – a goal that has been rarely explored. To address this question, we systematically analyze the energy distribution differences of clean, adversarial, and generated samples across various JEM variants and adversarially trained models. We observe that AT tends to reduce the energy gap between clean and adversarial samples, while JEMs reduce the gap between clean and synthetic ones. This observation suggests a key insight: if the energy distributions of all three data types can be aligned, we might unify the strengths of AT and JEMs, resolving their inherent trade-offs. Building on this idea, we propose Energy-based Joint Distribution Adversarial Training (EB-JDAT), to jointly model the clean data distribution, the adversarial distribution, and the classifier by maximizing their joint probability. To capture the full adversarial distribution, we introduce a new min-max energy optimization strategy. EB-JDAT is a general and flexible optimization method, compatible with various JEM variants. Extensive experimental results demonstrate that EB-JDAT not only maintains near original accuracy and generative capability of JEMs, but also significantly enhances robustness, even surpassing state-of-the-art ATs.
@@ 1 Introduction
Recently, Joint Energy-based Models (JEMs) grathwohl2019your ; yang2021jem++ ; yang2023towards have garnered significant attention for their ability to integrate generative modeling within a discriminative framework. This approach effectively bridges classification and generation from an energy-based perspective and exhibits surprising inherent robustness. However, JEMs still exhibit limited robustness compared to Adversarial Training (AT) mkadry2017towards . On the other hand, AT jia2022adversarial ; hammoudeh2024provable ; bortolussi2024robustness ; fang20253sat is widely recognized as the most effective method for improving the robustness of discriminative classifiers. However, it typically sacrifices accuracy on clean data and lacks generative capability zhang2019theoretically . Consequently, the inherent gap in classification, robustness and generation raises a natural question: Can a single model simultaneously achieve high classification accuracy, adversarial robustness, and generative performance? – a goal that has been rarely explored.
@@
Table 1: Mean and variance of one-to-one energy differences between clean and adv.
@@
To answer this question, we begin by investigating, from an energy perspective, the underlying causes of the performance gap between AT mkadry2017towards ; zhang2019theoretically and JEMs grathwohl2019your ; yang2023towards in terms of accuracy, generative capability, and adversarial robustness. We compute the energy distributions of clean vs. adversarial vs. generated samples by utilizing different models, and the results are shown in Fig. 1. First, the energy distributions of clean vs. adversarial samples, give a clear indication of AT tends to reduce the energy gap between clean and adversarial samples. Interestingly, JEMs implicitly reduce the gap between them, even though the distribution of adversarial samples is not explicitly incorporated during training, which explains the potential source of robustness in JEMs. Further, we compute the mean and variance of one-to-one energy differences between them. As reported in Tab. 1, models with smaller energy differences exhibit higher robustness. Next, we investigate the classification and generation gaps by analyzing the energy distributions of clean vs. generated samples in Fig. 1. Compared to AT, JEMs reduce the gap between them, achieving high accuracy in classification while maintaining strong generative quality. Building on these observations, we draw a key insight: if the energy distributions of all three data types can be aligned, we might unify the strengths of AT and JEMs, resolving their inherent trilemma.
Inspired by this idea, we align the energy distribution of the three data types by jointly modeling the clean data distribution, the adversarial distribution and the classifier, by learning $p_{\theta}(\mathbf{x},\tilde{\mathbf{x}},y)$ via energy-based model, where $\mathbf{x}$, $y$ and $\tilde{\mathbf{x}}$ denote the clean samples, labels and adversarial examples, and $\theta$ denote the model parameters. Through Bayesian decomposition, $p_{\theta}(\mathbf{x},\tilde{\mathbf{x}},y)$ can be further factorized as the data distribution $p_{\theta}(\mathbf{x})$, adversarial distribution $p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})$ and $p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})$. While $p_{\theta}(\mathbf{x})$ can be computed via approximated sampling methods nijkamp2020anatomy and $p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})$ is simply a cross-entropy objective for robust classification, computing $p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})$ is not straightforward because the full adversarial distribution is not observed during the training phase. However, our key observation is that adversarial perturbations almost surely move samples out of the original data manifold, which lies in high-density areas, into low-density areas where the classifier is more likely to fail. Motivated by this observation, we propose a new min-max energy optimization to approximate $p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})$. During the sampling phase, we search for high-energy adversarial examples, then minimize the energy gap between adversarial and clean samples during the training phase, thereby pulling the adversarial samples back into low-energy regions. We name our proposed method as Energy-based Joint Distribution Adversarial Training, or EB-JDAT. Extensive experiments across multiple datasets and attacks show that EB-JDAT achieves 64.76% and 35.63% robustness (under AutoAttack) on CIFAR-10 and CIFAR-100 datasets, outperforming the SOTA AT methods by +9.24% and +4.86% respectively, while maintaining near original clean accuracy and generative capability, which is competitive to most advanced JEMs.
@@ 2 Related work
Since the discovery of adversarial vulnerabilities in deep learning models yuan2021automa ; ma2023transferable , numerous defense strategies have been proposed to mitigate this security challenge. Prominent approaches include input denoising li2022feature , gradient regularization li2023revisiting ; jia2022prior , defensive distillation huang2023boosting , and adversarial training (AT) mkadry2017towards ; fang20253sat . Recent studies confirm that AT remains among the most effective defense methods sengphanith2023evaluating , significantly enhancing model robustness by incorporating adversarial examples into the training process. Notably, mkadry2017towards first formulated AT as a saddle-point optimization problem (PGD), laying the foundational framework for subsequent robust training methodologies. As adversarial attack methods have evolved, various novel strategies have emerged, refining conventional AT methods jia2022adversarial ; hammoudeh2024provable ; bortolussi2024robustness ; fang20253sat . However, many of these strategies prioritize robustness at the expense of clean accuracy. which is the principal evaluation metric in visual classification, highlighting an inherent trade-off in AT. To mitigate this issue, zhang2019theoretically proposed TRADES, which constrains the divergence in output probability between clean and adversarial examples. Additionally, co2022jacobian combined Jacobian regularization with model ensembling, introducing the Jacobians ensemble approach that effectively balances robustness and accuracy against universal adversarial perturbations. Furthermore, DUCAT wang2024new assigns a dummy class to each original class to absorb hard adversarial examples during training and remaps them back at inference to mitigating the trade-off. Although these approaches represent notable advancements, they still suffer from substantial accuracy degradation relative to standard classifiers, while offering only modest enhancements in robustness. Therefore, achieving a trade-off between accuracy and robustness remains an unresolved challenge in AT.
@@
Recent studies have indicated that features extracted by robust classifiers are more closely aligned with human perceptual mechanisms compared to those derived from standard classifiers wangunified . The Joint Energy-based Model (JEM) grathwohl2019your reformulates conventional softmax classifiers into an energy-based model (EBM), facilitating hybrid discriminative-generative modeling and clarifying the connection between robustness and generative modeling. More recent efforts have focused on improving training efficiency and speed in JEM yang2021jem++ ; yang2023towards , effectively bridging the gap between classification and generation. Additionally, several approaches specifically focus on enhancing robustness, from an energy perspective. JEAT zhu2021towards initially demonstrated a link between AT and EBM, highlighting that despite differences in how these methods modify the energy function, they ultimately employ comparable contrastive strategies. Subsequently, mirza2024shedding reinterpreted robust classifiers within an energy-based framework and analyzed the energy landscape to reveal differential effects of targeted and non-targeted attacks during AT. However, these methods often sacrifice robustness in exchange for notable reductions in accuracy and generative capacity, with some even proving less robust than traditional adversarial defense strategies. Thus, bridging the inherent gap in classification, generation, and robustness is challenging, yet no existing method has effectively addressed this challenge. Although SOTA JEM variants perform well in classification and generative capabilities, their robustness remains limited.
@@ 3.1 Preliminaries
Energy-based Models (EBMs)
From an energy perspective, the clean data distribution ${p}_{\theta}(\mathbf{x})$ can be explicitly parameterized through an energy function grathwohl2019your :
@@
where $E_{\theta}(\cdot):\mathbb{R}^{D}\rightarrow\mathbb{R}$ denotes the energy function. This energy function assigns each data point a scalar value representing its density within the continuous data distribution. The normalizing constant $Z(\theta)=\int_{\mathbf{x}}\exp(-E_{\theta}(\mathbf{x}))$ ensures the proper normalization of probabilities. During training, the energy function $E_{\theta}(\mathbf{x})$ is optimized to assign lower energy values to high-density regions (e.g., training data) and higher energy values to low-density regions (e.g., adversarial samples). To optimize the energy-based model, a common strategy is to perform maximum likelihood estimation of the parameters $\theta$, in which the gradient of the log-likelihood is expressed as follows:
$\frac{\partial{log}p_{\theta}(\mathbf{x})}{\partial\theta}=\mathbb{E}_{p_{\theta}(\mathbf{x^{\prime}})}\left[\frac{\partial E_{\theta}(\mathbf{x^{\prime}})}{\partial\theta}\right]-\frac{\partial E_{\theta}(\mathbf{x})}{\partial\theta},$ (2)
where $\mathbf{x^{\prime}}$ is simply a sample obtained by sampling from the model distribution.
@@
JEM grathwohl2019your reinterprets the standard softmax classifier within the framework of an EBM. Specifically, it redefines the logits of the classifier $f_{\theta}$ to model the joint distribution $p_{\theta}(\textbf{x},y)$ of the input data x and their labels $y$ using an energy-based approach:
@@
In JEMs grathwohl2019your ; yang2021jem++ ; yang2023towards , the model parameters $\theta$ are optimized by maximizing $\log p_{\theta}(\mathbf{x},y)$:
@@
Adversarial Training (AT)
Adversarial training can be viewed as a min-max optimization problem that integrates inner loss maximization and outer loss minimization mkadry2017towards . Given data $\mathbf{x}\in\mathbf{X}$, its corresponding label $y\in\mathbf{y}$, and a classification model $f_{\theta}$, the objective of the inner maximization is to identify the most aggressive adversarial examples by maximizing the loss. Subsequently, the outer minimization seeks to minimize the average loss of these adversarial examples. The specific optimization process is detailed in:
@@ 3.2 Motivation
ATmkadry2017towards has demonstrated strong robustness against adversarial attacks. However, the adversarial samples introduced during training lie outside the true data manifold, causing the model to overemphasize these adversarial features while neglecting meaningful patterns in clean data ilyas2019adversarial . Consequently, AT-trained models exhibit a noticeable accuracy gap compared to standard classifiers. In contrast, JEMs have shown the ability to reinterpret discriminative classifiers as generative models grathwohl2019your , achieving notable performance in both image classification and generation. While JEMs exhibit improved adversarial robustness relative to standard models, a significant robustness gap still remains when compared to AT-trained methods. This naturally raises a new research question: Can we combine the strengths of AT and JEMs to develop a unified framework that mitigates, or even eliminates, triple trade-off between accuracy, robustness and generative performance?
To answer this question, we begin by visualizing the energy distributions of clean vs. adversarial vs. generated samples to investigate the source of the performance gains in JEMs and AT, which are shown in Fig. 1. We first analyze the energy distributions of clean vs. adversarial samples. Obviously, for models trained by AT, the energy distributions of clean and adversarial samples almost completely overlap. While for JEMs, the two distributions partially overlap but less than that observed in AT-based methods. Further, We compute the mean and variance of one-to-one energy differences between clean samples and their corresponding adversarial examples and report the results in Tab. 1. By analyzing these results, we have two findings: 1) the greater the overlap between the energy distributions of clean and adversarial samples, the higher the model’s robustness. 2) JEMs implicitly reduce the gap between adversarial and clean data, although the distribution of adversarial samples is not explicitly incorporated during training. This also explains the source of robustness in JEMs.
@@
3.3 Energy-based Joint Adversarial Distribution Training
@@
$p_{\theta}(\mathbf{x},\tilde{\mathbf{x}},y)=p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})p_{\theta}(\tilde{\mathbf{x}},\mathbf{x})=p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})p_{\theta}(\mathbf{x}),$ (6)
where $\theta$ denotes the model parameters. The conditional distribution ${p}_{\theta}\left({y}|\tilde{\mathbf{x}},{\mathbf{x}}\right)$ denotes cross-entropy objective for robust classification, such as AT mkadry2017towards . $p_{\theta}(\tilde{\mathbf{x}},\mathbf{x})$ aims to capture the joint distribution of clean and adversarial examples, and it can be further decomposed using Bayesian theorem as $p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})p_{\theta}(\mathbf{x})$. Similar to Eq. 1, the adversarial distribution $p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})$ can also be parameterized through an energy-based formulation:
$p_{\theta}(\tilde{\mathbf{x}}|\mathbf{x})=\frac{\exp\left(-E_{\theta}(\tilde{\mathbf{x}}|\mathbf{x})\right)}{\tilde{Z}_{\theta}}=\frac{\sum_{y\in\mathbf{y}}\exp\left(f_{\theta}(\tilde{\mathbf{x}})[y]\right)}{\tilde{Z}_{\theta}},$ (7)
where $\tilde{Z}(\theta)=\int_{\tilde{\mathbf{x}}}\exp(-E_{\theta}(\tilde{\mathbf{x}}))$. A natural choice to optimize the model parameters $\theta$ is to maximize the log-likelihood of $p_{\theta}(\mathbf{x},\tilde{\mathbf{x}},y)$:
$log\ p_{\theta}(\mathbf{x},\tilde{\mathbf{x}},y)=log\ p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})+log\ p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})+log\ p_{\theta}(\mathbf{x}),$ (8)
where $log\ p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})$ is simply a cross-entropy objective for robust classifier, $log\ p_{\theta}(\mathbf{x})$ can be estimated via various sampling methods welling2011bayesian . In our method, Eq. 2 is approximated by nijkamp2020anatomy :
$\frac{\partial{log}p_{\theta}(\mathbf{x})}{\partial\theta}\approx\frac{\partial}{\partial\theta}\left[\frac{1}{L_{1}}\sum_{i=1}^{L_{1}}E_{\theta}\left(\mathbf{x}_{i}^{+}\right)\right.\\ \left.-\frac{1}{L_{2}}\sum_{i=1}^{L_{2}}E_{\theta}\left(\mathbf{x}_{i}^{-}\right)\right],$ (9)
where $\left\{\mathbf{x}_{i}^{+}\right\}_{i=1}^{L_{1}}$ denote all the training samples in a batch, and $\left\{\mathbf{x}_{i}^{-}\right\}_{i=1}^{L_{2}}$ are i.i.d samples drawn from $p_{\theta}(\mathbf{x})$ via Stochastic Gradient Langevin Dynamics (SGLD) welling2011bayesian :
@@
where $c$ is a step size, $\mathbf{N}$ denotes the normal distribution, and $\mathbf{I}$ is the identity matrix. Therefore, the key to optimizing Eq. 8 is to estimate $log\ p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})$. Although the full adversarial distribution is not observed during training, we start from a straightforward yet key observation from an energy view: adversarial samples typically appear near the boundaries or even outside the support of the original data distribution, situating them in low-density, thus high-energy regions. Additionally, by visualizing the energy distributions of clean and adversarial samples in Fig. 1 and Tab. 1, we observe that, under robust classifiers, these distributions are nearly indistinguishable. This suggests that, after AT, adversarial examples are pulled back from high-energy area into low-energy (i.e., high-density) areas, close to clean samples. This inspires us to propose a min-max optimization framework to compute $log\ p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})$:
@@
$\min_{\theta}\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}}\left[\max_{\tilde{\mathbf{x}}}\left(E_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})-E_{\theta}(\mathbf{x})\right)\right]\;\text{s.t.}\;\tilde{\mathbf{x}}\;\text{is natural}\;,$ (11)

### v2

@@ Abstract
Joint Energy-based Models (JEMs) are well known for their ability to unify classification and generation within a single framework. Despite their promising generative and discriminative performance, their robustness remains far inferior to adversarial training (AT), which, conversely, achieves strong robustness but sacrifices clean accuracy and lacks generative ability. This inherent trilemma—balancing classification accuracy, robustness, and generative capability—raises a fundamental question: Can a single model achieve all three simultaneously? To answer this, we conduct a systematic energy landscape analysis of clean, adversarial, and generated samples across various JEM and AT variants. We observe that AT reduces the energy gap between clean and adversarial samples, while JEMs narrow the gap between clean and synthetic ones. This observation suggests a key insight: if the energy distributions of all three data types can be aligned, we might bridge their performance disparities. Building on this idea, we propose Energy-based Joint Distribution Adversarial Training (EB-JDAT), a unified generative-discriminative-robust framework that maximizes the joint probability of clean and adversarial distribution. EB-JDAT introduces a novel min–max energy optimization to explicitly aligning energies across clean, adversarial, and generated samples. Extensive experiments on CIFAR-10, CIFAR-100, and ImageNet subsets demonstrate that EB-JDAT achieves state-of-the-art robustness while maintaining near-original accuracy and generation quality of JEMs, effectively resolving the triple trade-off between accuracy, robustness, and generation.
@@ 1 Introduction
Figure 1: Comparisons of SOTA AT-based methods on CIFAR-10 in terms of accuracy and robustness (AutoAttack). Our method achieves the best robustness while maintaining competitive standard accuracy.
Recently, Joint Energy-based Models (JEMs) [13, 39, 41] have garnered significant attention for their ability to integrate generative modeling within a discriminative framework. This approach effectively bridges classification and generation from an energy-based perspective and exhibits surprising inherent robustness. However, JEMs still exhibit limited robustness compared to Adversarial Training (AT) [25]. On the other hand, AT [19, 14, 2, 45] is widely recognized as the most effective method for improving the robustness of discriminative classifiers. However, it typically sacrifices accuracy on clean data and lacks generative capability [44]. Consequently, the inherent gap in classification, robustness and generation raises a natural question: Can a single model simultaneously achieve high classification accuracy, adversarial robustness, and generative performance? – a goal that has been rarely explored.
@@
To answer this question, we investigate the underlying causes of the performance gap between AT [25, 44] and JEMs [13, 41] in terms of accuracy, generative capability, and adversarial robustness from an energy perspective. We compute the energy distributions of clean, adversarial, and generated samples, with results shown in Fig. 2. Further, we compute the mean and variance of the one-to-one energy differences between clean and adversarial samples, as reported in Tab. 1. The energy distributions shows that AT reduces the energy gap between clean and adversarial samples, thereby bringing robustness. In contrast, JEMs reduce the gap between clean and generated samples, which brings generative capability and higher accuracy. Based on these insights, we draw a key conclusion: if the energy distributions of all three data types can be aligned, we might unify the strengths of AT and JEMs, resolving their inherent trilemma.
Table 1: Mean and variance of one‐to‐one energy differences between clean and adv. Results in bold indicate the best.
@@
Inspired by this idea, we align the energy distributions of clean, adversarial and generated data through a joint energy-based model of $p_{\theta}(\mathbf{x},\tilde{\mathbf{x}},y)$, with $\mathbf{x}$ clean samples, $\tilde{\mathbf{x}}$ adversarial examples, $y$ class labels and $\theta$ the model parameters. Through Bayesian decomposition, $p_{\theta}(\mathbf{x},\tilde{\mathbf{x}},y)$ can be further factorized as the data distribution $p_{\theta}(\mathbf{x})$, adversarial distribution $p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})$ and $p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})$. While $p_{\theta}(\mathbf{x})$ can be computed via approximated sampling methods [27] and $p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})$ is simply a cross-entropy objective for robust classification, computing $p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})$ is not straightforward because the full adversarial distribution is not observed during the training phase. However, our key observation is that adversarial perturbations almost surely move samples out of the original data manifold, which lies in high-density areas, into low-density areas where the classifier is more likely to fail, as illustrated by the weakly robust classifiers JEM [13] and SADAJEM [41] in Fig. 2. Motivated by this observation, we propose a new min-max energy optimization to approximate $p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})$. During the sampling phase, we search for high-energy adversarial examples, then minimize the energy gap between adversarial and clean samples during the training phase, thereby pulling the adversarial samples back into low-energy regions. We name our proposed method as Energy-based Joint Distribution Adversarial Training, or EB-JDAT. Extensive experiments across multiple datasets and attacks show that EB-JDAT achieves 68.76%, 35.63% and 32.40% robustness (under AutoAttack [6]) on CIFAR-10, CIFAR-100 and ImageNet subset, outperforming the SOTA AT methods by +10.78%, +4.70% and +7.88% respectively, while maintaining near original clean accuracy and generative capability, which is competitive to most advanced JEMs.
@@ 2 Related Work
Since the discovery of adversarial vulnerabilities in deep learning models [42, 24], numerous defense strategies have been proposed to mitigate this security challenge. Prominent approaches include input denoising [21], gradient regularization [22, 18], defensive distillation [16], and adversarial training (AT) [25, 10, 45]. Recent studies confirm that AT remains among the most effective defense methods [29], significantly enhancing model robustness by incorporating adversarial examples into the training process. Notably, [25] first formulated AT as a saddle-point optimization problem (PGD), laying the foundational framework for subsequent robust training methodologies. As adversarial attack methods have evolved, various novel strategies have emerged, refining conventional AT methods [19, 14, 2, 10, 45]. However, many of these strategies prioritize robustness at the expense of clean accuracy. which is the principal evaluation metric in visual classification, highlighting an inherent trade-off in AT. To mitigate this issue, [44] proposed TRADES, which constrains the divergence in output probability between clean and adversarial examples. Additionally, [4] combined Jacobian regularization with model ensembling, introducing the Jacobians ensemble approach that effectively balances robustness and accuracy against universal adversarial perturbations. Furthermore, DUCAT [32] assigns a dummy class to each original class to absorb hard adversarial examples during training and remaps them back at inference to mitigating the trade-off. Although these approaches represent notable advancements, they still suffer from substantial accuracy degradation relative to standard classifiers, while offering only modest enhancements in robustness. Therefore, achieving a trade-off between accuracy and robustness remains an unresolved challenge in AT.
@@
Recent studies have indicated that features extracted by robust classifiers are more closely aligned with human perceptual mechanisms compared to those derived from standard classifiers [33]. The Joint Energy-based Model (JEM) [13] reformulates conventional softmax classifiers into an energy-based model (EBM), facilitating hybrid discriminative-generative modeling and clarifying the connection between robustness and generative modeling. More recent efforts have focused on improving training efficiency and speed in JEM [39, 41], effectively bridging the gap between classification and generation. Additionally, several approaches specifically focus on enhancing robustness, from an energy perspective. JEAT [46] initially demonstrated a link between AT and EBM, highlighting that despite differences in how these methods modify the energy function, they ultimately employ comparable contrastive strategies. Subsequently, [26] reinterpreted robust classifiers within an energy-based framework and analyzed the energy landscape to reveal differential effects of targeted and non-targeted attacks during AT. However, these methods often sacrifice robustness in exchange for notable reductions in accuracy and generative capacity, with some even proving less robust than traditional adversarial defense strategies. Thus, bridging the inherent gap in classification, generation, and robustness is challenging, yet no existing method has effectively addressed this challenge. Although SOTA JEM variants perform well in classification and generative capabilities, their robustness remains limited.
@@ 3.1 Preliminaries
Energy-based Models (EBMs).
From an energy perspective, the clean data distribution ${p}_{\theta}(\mathbf{x})$ can be explicitly parameterized through an energy function [13]:
@@
where $E_{\theta}(\cdot):\mathbb{R}^{D}\rightarrow\mathbb{R}$ denotes the energy function, $f_{\theta}(x)[y]$ denotes the $y^{\mathrm{th}}$ index of $f_{\theta}[y]$, i.e., the logit corresponding the $y^{\mathrm{th}}$ class label. This energy function assigns each data point a scalar value representing its density within the continuous data distribution. The normalizing constant $Z(\theta)=\int_{\mathbf{x}}\exp(-E_{\theta}(\mathbf{x}))$ ensures the proper normalization of probabilities. During training, the energy function $E_{\theta}(\mathbf{x})$ is optimized to assign lower energy values to high-density regions (e.g., training data) and higher energy values to low-density regions (e.g., adversarial samples). To optimize the energy-based model, a common strategy is to perform maximum likelihood estimation of the parameters $\theta$, in which the gradient of the log-likelihood is expressed as follows:
$\frac{\log p_{\theta}(\mathbf{x})}{\partial\theta}=\mathbb{E}_{p_{\theta}(\mathbf{x^{\prime}})}\left[\frac{\partial E_{\theta}(\mathbf{x^{\prime}})}{\partial\theta}\right]-\frac{\partial E_{\theta}(\mathbf{x})}{\partial\theta},$ (2)
where $\mathbf{x^{\prime}}$ is simply a sample obtained by sampling the distribution of models.
@@
JEM [13] reinterprets the standard softmax classifier within the framework of an EBM. Specifically, it redefines the logits of the classifier $f_{\theta}$ to model the joint distribution $p_{\theta}(\textbf{x},y)$ of the input data x and their labels $y$ using an energy-based approach:
@@
In JEMs [13, 39, 41], the model parameters $\theta$ are optimized by maximizing $\log p_{\theta}(\mathbf{x},y)$:
@@
Adversarial Training (AT).
Adversarial training can be viewed as a min-max optimization problem that integrates inner loss maximization and outer loss minimization [25]. Given data $\mathbf{x}\in\mathbf{X}$, its corresponding label $y\in\mathbf{y}$, and a classification model $f_{\theta}$, the objective of the inner maximization is to identify the most aggressive adversarial examples by maximizing the loss. Subsequently, the outer minimization seeks to minimize the average loss of these adversarial examples. The specific optimization process is detailed in:
@@ 3.2 Motivation
AT[25] has demonstrated strong robustness against adversarial attacks. However, the adversarial samples introduced during training lie outside the true data manifold, causing the model to overemphasize these adversarial features while neglecting meaningful patterns in clean data [17]. Consequently, AT-trained models exhibit a noticeable accuracy gap compared to standard classifiers. In contrast, JEMs have shown the ability to reinterpret discriminative classifiers as generative models [13], achieving notable performance in both image classification and generation. While JEMs exhibit improved adversarial robustness relative to standard models, a significant robustness gap still remains when compared to AT-trained methods. This naturally raises a new research question: Can we combine the strengths of AT and JEMs to develop a unified framework that mitigates, or even eliminates, triple trade-off between accuracy, robustness and generative performance? To answer this question, we begin by visualizing the energy distributions of clean vs. adversarial vs. generated samples to investigate the source of the performance gains in JEMs and AT, which are shown in Fig. 2.
Analysis of the Robustness.
We first analyze the energy distributions of clean vs. adversarial samples. Obviously, for models trained by AT, the energy distributions of clean and adversarial samples almost completely overlap. While for JEMs, the two distributions partially overlap but less than that observed in AT-based methods. Further, We compute the mean and variance of one-to-one energy differences between clean samples and their corresponding adversarial examples and report the results in Tab. 1. By analyzing these results, we have two findings: 1) the greater the overlap between the energy distributions of clean and adversarial samples, the higher the model’s robustness. 2) JEMs implicitly reduce the gap between adversarial and clean data, although the distribution of adversarial samples is not explicitly incorporated during training. This also explains the source of robustness in JEMs.
Analysis of the Generative Capability.
@@
3.3 Energy-based Joint Distribution Adversarial Training
@@
$\begin{split}p_{\theta}(\mathbf{x},\tilde{\mathbf{x}},y)&=p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})\,p_{\theta}(\tilde{\mathbf{x}},\mathbf{x})\\ &=p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})\,p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})\,p_{\theta}(\mathbf{x}),\end{split}$ (6)
where $\theta$ denotes the model parameters. The conditional distribution ${p}_{\theta}\left({y}|\tilde{\mathbf{x}},{\mathbf{x}}\right)$ denotes cross-entropy objective for robust classification, such as AT [25]. $p_{\theta}(\tilde{\mathbf{x}},\mathbf{x})$ aims to capture the joint distribution of clean and adversarial examples, and it can be further decomposed using Bayesian theorem as $p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})p_{\theta}(\mathbf{x})$. Similar to Eq. 1, the adversarial distribution $p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})$ can also be parameterized through an energy-based formulation:
$p_{\theta}(\tilde{\mathbf{x}}|\mathbf{x})=\frac{\exp\left(-E_{\theta}(\tilde{\mathbf{x}}|\mathbf{x})\right)}{\tilde{Z}_{\theta}}=\frac{\sum_{y\in\mathbf{y}}\exp\left(f_{\theta}(\tilde{\mathbf{x}}|\mathbf{x})[y]\right)}{\tilde{Z}_{\theta}},$ (7)
where $\tilde{Z}(\theta)=\int_{\tilde{\mathbf{x}}}\exp(-E_{\theta}(\tilde{\mathbf{x}}|\mathbf{x}))$. A natural choice to optimize the model parameters $\theta$ is to maximize the log-likelihood of $p_{\theta}(\mathbf{x},\tilde{\mathbf{x}},y)$:
$\log p_{\theta}(\mathbf{x},\tilde{\mathbf{x}},y)$ $=\log p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})$
$\quad+\log p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})+\log p_{\theta}(\mathbf{x}),$ (8)
where $\log p_{\theta}(y\mid\tilde{\mathbf{x}},\mathbf{x})$ is simply a cross-entropy objective for robust classifier, $\log p_{\theta}(\mathbf{x})$ can be estimated via various sampling methods [36]. In our method, Eq. 2 is approximated by [27]:
$\frac{\partial{\log}p_{\theta}(\mathbf{x})}{\partial\theta}\approx\frac{\partial}{\partial\theta}\left[\frac{1}{L_{1}}\sum_{i=1}^{L_{1}}E_{\theta}\left(\mathbf{x}_{i}^{+}\right)\right.\\ \left.-\frac{1}{L_{2}}\sum_{i=1}^{L_{2}}E_{\theta}\left(\mathbf{x}_{i}^{-}\right)\right],$ (9)
where $\left\{\mathbf{x}_{i}^{+}\right\}_{i=1}^{L_{1}}$ denote all the training samples in a batch, and $\left\{\mathbf{x}_{i}^{-}\right\}_{i=1}^{L_{2}}$ are i.i.d samples drawn from $p_{\theta}(\mathbf{x})$ via Stochastic Gradient Langevin Dynamics (SGLD) [36]:
@@
where $c$ is a step size, $\mathbf{N}$ denotes the normal distribution, and $\mathbf{I}$ is the identity matrix. Therefore, the key to optimizing Sec. 3.3 is to estimate $\log\ p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})$. Although the full adversarial distribution is not observed during training, we start from a straightforward yet key observation from an energy view: adversarial samples typically appear near the boundaries or even outside the support of the original data distribution, situating them in low-density, thus high-energy regions. Additionally, by visualizing the energy distributions of clean and adversarial samples in Fig. 2 and Tab. 1, we observe that, under robust classifiers, these distributions are nearly indistinguishable. This suggests that, after AT, adversarial examples are pulled back from high-energy area into low-energy (i.e., high-density) areas, close to clean samples. This inspires us to propose a min-max optimization framework to compute $\log p_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})$:
@@
$\min_{\theta}\mathbb{E}_{(\mathbf{x},y)\sim\mathcal{D}}\left[\max_{\|\tilde{\mathbf{x}}-\mathbf{x}\|\in\Omega}\left(E_{\theta}(\tilde{\mathbf{x}}\mid\mathbf{x})-E_{\theta}(\mathbf{x})\right)\right],$ (11)